Status Report

NASA Space Station On-Orbit Status 28 July 2004

By SpaceRef Editor
July 28, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 28 July 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally except those noted previously or below.  

EVA-10 preparations are continuing throughout the day, starting with leak checks and valve functionality tests on the suits and their BSS interface units, supported by tagup with ground specialists.

Today’s pre-EVA ops also include pressure checks of the primary & backup oxygen tanks (BK-3), and function tests of the hatch KVDs (pressure equalization valves, PEVs) via the POVs (EVA support panels), both in the Service Module Transfer Compartment (SM PkhO) and in the DC1 docking compartment.  This is to back up an earlier checkout of these vital elements.

Further, the crew will test the Russian BETA-08 ECG (electrocardiogram) lead cable belts, worn under the Orlan-M suits, for their function, using the Gamma-1M medical complex.

After setting up communications links, the crew performs checks on Orlan and BSS telemetry, voice, and biomedical parameter transmission.

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

CDR Padalka terminated the bake-out cycle on the BMP micropurification system’s channel 2, moding the channel back to Purify.  This restored both filter beds to Purify mode.  [The regeneration of the air purifier filter beds is repeated every 20 days.  Each bakeout to space vacuum takes about 24 hours.]  — (This was wrongly reported for yesterday, instead of the CDR’s switching to filter #2 regen).

Preparatory to the Progress 15P docking on 8/14, both crewmembers conducted a test of the TORU teleoperator control system between the SM and the docked Progress 14P.  Progress thrusters (DPO) were inhibited and not involved.  The test was performed over RGS (Russian ground sites) on Daily Orbits 3.   [TORU is the manual mode through which Padalka can perform necessary guidance functions from the SM in the event of a failure of the “Kurs” automated rendezvous and docking (AR&D) of the Progress.  He would control the cargo ship’s motions from a control panel, viewing the approach to the ISS via the Klest-M television camera mounted on the Progress.  Remote TORU control from the ground is not available.  The test was scheduled for today since 14P will be gone after 7/30.]

Later in the day, the crew finished up loading of the Progress with excessed equipment and trash.  Then, Gennady Padalka, assisted by the FE as required, removed no-longer-needed electronic equipment from the cargo ship.   [After Central Computer control was transferred from 14P thrusters to the Service Module (SM) yaw, pitch and roll thrusters, the US-21 matching unit and SKV-1 dehumidifier were deactivated, followed by disconnecting the cables of the BITS 2-12 onboard telemetry measurement system and turning off its VD-SU monitoring mode.  The crew then unbolted and removed the Progress’ US-21 matching unit in its container box.  BITS was later reactivated, as was SKV-1.  The US-21, with its associated commutator gear, provides the electronic interface between the SM and the Progress for SM computer control of the Progress thrusters.  With 14P being readied for separation and jettisoning, the valuable electronics were retained in storage, to be returned to Earth and recycled on a future Progress.]

FE/SO Mike Fincke conducted the monthly PEP (portable emergency provisions) audit and inspection.   [The procedure involves verification that PFEs (portable fire extinguishers), PBAs (portable breathing assemblies), QDMAs (quick-don mask assemblies) and EHTKs (extension hose/tee kits) are free of damage to ensure their functionality, and to track shelf life/life cycles on the hardware.  PEPs are not removed from their locker unless obvious damage is discovered during the inspection.  There are a total of 5 PBAs in the U.S. segment (USOS), viz., two in the Node, two in the Lab, and one in the Airlock.]

Mike also completed the periodic one-hour inspection of emergency lighting power sources (ELPS) in the Node.   [There are three ELPSs in “Unity”, two in the Lab “Destiny”, and one ELPS in the Airlock “Quest”.]

Regular daily routine tasks were performed by Mike (Lab payload status checkup) and Gennady (SOZh life support system maintenance & inventory management system [IMS] delta file preparation).

At 1:30pm, the FE is scheduled to transfer coolant water from the MTL (medium temperature loop) to the LTL (low temperature loop) of the U.S. ITCS (internal thermal control system). [Every time the crew performs the mode transitions and necessary “jumpering” (hose connecting) to place the ITCS in the unmanned configuration for EVA, a slight amount of water is transferred from the LTL to the MTL.  As a result, the MT accumulator quantity has been bumped up against the level associated with an ammonia leak warning setting (90%). Rather than inhibiting the warning Message and associated FDIR at this level, Mike will move some of that water back into the LT Loop.]

Fincke collected the previously (7/26) deployed passive U.S. FMK (formaldehyde monitoring kit) sampling badges from the Lab and SM for packing and return to Earth.

Yesterday’s repressurization of the station with 14 mmHg (Torr) of oxygen (O2) from Progress tankage brought total cabin pressure to 754 mmHg.  The MCA (Mass Constituent Analyzer) was turned on, but the zero calibration was not performed until after the repress.  The MCA ppO2 reading after repress was 177.22 mmHg (approximately 23.5% of the atmosphere).  A second O2 repress adding another 3-4 mmHg is being performed today since 5:30am EDT.  [The ECLS team evaluated recent MCA calibration data and determined that the MCA error band could be reduced from 6 mmHg to 2 mmHg for the two Progress 14P represses.]

The pump swap from the ‘A’ pump to the ‘B’ pump of loop A of the EETCS PFCS (Early External Thermal Control System/Pump Flow Control Subassembly) was successfully accomplished yesterday. The EETCS PFCS loop B pumps will be swapped today.  [The ‘B’ pumps in both loops will run for a year and then be swapped back to the ‘A’s again.  The ‘A’ pumps were activated during Mission 4A.]

Also yesterday, the ground enabled the RGA FDIR (Rate Gyro Assembly/Fault Detection, Isolation and Recovery) with a new PPL (Pre-Positioned Load).  The new software allows the software to determine whether or not the laser rate gyros are providing false data.  [This was previously inhibited due to repeated false data flags.  GN&C (guidance, navigation & control) specialists have been reviewing flight data since the beginning of April, and are satisfied that the incorporation of the updates will allow the FDIR to operate correctly.]

The crew performed their regular daily physical exercise on TVIS treadmill, RED resistive exerciser, and VELO stationary bike with load trainer.

Major upcoming events:

  • EVA procedures review, 14P undock prep — 7/29;
  • Orlan training run, battery charging, camera prep — 7/30;
  • Progress 14P undocking — 7/30 (2:06am EDT);
  • EVA timeline review — 8/1;
  • OpsLan reconfig, etc. — 8/2;
  • Orlan EVA-10 from DC-1 — 8/3 (hatch open: 2:50am EDT);
  • EVA debrief, etc. — 8/4;
  • Progress 15P launch — 8/11 (1:01am EDT);
  • Progress 15P docking — 8/14 (2:05am EDT).

Today’s CEO photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Hawaii (looking left two degrees for the Big Island),Darwin, Australia(nadir pass),Sydney, Australia(nadir pass),Aurora, Australia(looking right.  Auroral activity varies daily, and has been very strong),Persian Gulf(ISS approached the high contrast city/coast region on a descending pass following the north coast of the gulf: looking right to the cities along the south coast),and Muscat, Oman(nadir pass).

Mike Fincke’s current (7/27) Crew Earth Observation (CEO) image count has reached a healthy 10,275. 

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 9 crew visit:

Expedition 9 Flight Crew Plans can be found at

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

  • ISS Orbit  (as of this morning, 2:47am EDT [= epoch]):
  • Mean altitude — 357.9 km
  • Apogee height — 362.1 km
  • Perigee height — 353.7km
  • Period — 91.7 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0006282
  • Solar Beta Angle — 2.8 deg
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 100 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 32477

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.